# Value Residual Learning For Alleviating Attention Concentration In Transformers

### v2

@@ Abstract
Transformers can capture long-range dependencies using self-attention, allowing tokens to attend to all others directly. However, stacking multiple attention layers leads to attention concentration. One natural way to address this issue is to use cross-layer attention, allowing information from earlier layers to be directly accessible to later layers. However, this approach is computationally expensive. To address this problem, we propose Transformer with residual value (ResFormer) which approximates cross-layer attention through adding a residual connection from the values of the the first layer to all subsequent layers. Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer, reducing the $KV$ cache by nearly 50%. Comprehensive empirical evidence demonstrates that ResFormer mitigates attention concentration problem in deeper layers and enhances representation across most layers, outperforming the vanilla Transformer, DenseFormer, and NeuTRENO in training error as well as downstream tasks. Further visualization results suggest that Resformer alleviates attention sinks through avoiding value-state drains. SVFormer trains significantly faster than the vanilla Transformer and performs better than other methods like GQA and CLA, with performance influenced by sequence length and cumulative learning rate.  [^1] Code is available at https://github.com/Zcchill/Value-Residual-Learning.
@@ 1 Introduction
The Transformer (Vaswani, 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin, 2018, Lan, 2019, Brown, 2020) and computer vision tasks (Dosovitskiy, 2020). The discovery of scaling laws (Hoffmann et al., 2022, Kaplan et al., 2020) has driven the pursuit of larger Transformer models by increasing network depth and width.
@@
Existing solutions to alleviate the over-smoothing problem in Transformer include adding extra regularizers (Nguyen et al., 2023, Shi et al., 2022) and optimizing the information flow within the model (Pagliardini et al., 2024). During the era of convolutional neural network architectures, Stochastic Depth (Huang et al., 2016) reduces the likelihood of over-smoothing by randomly dropping layers during training and DenseNet (Huang et al., 2017) mitigates the impact of over-smoothing by allowing each layer to directly access the hidden states of all preceding layers. Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer. Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value states of the first layer and the current layer to the attention output.
@@
In this paper, we address the problem of multi-layer attention from a novel perspective. We introduce the phenomenon of attention concentration, which describes how a model’s attention increasingly focuses on fewer tokens as the network depth increases, as illustrated in Fig. 1 (Right). We quantify the degree of attention concentration using the entropy of the distribution of token importance, where lower entropy indicates a more pronounced concentration. In a deep model, the attention mechanism exhibits a pattern of “concentration - dispersion - concentration,” where the model could potentially lose useful information during the concentrated phase. See Fig. 18 for analysis of over-smoothing.
To alleviate the attention concentration from stacking multiple attention layers, an effective method is to use cross-layer attention on information from earlier layers. Given the high computational cost of cross-layer attention, we propose ResFormer as an approximation. ResFormer achieves a similar effect by applying a residual connection between the value states of the current layer and the first layer before the attention operation. Experimental results show that ResFormer mitigates the attention concentration effect in deeper layers, maintaining attention dispersion throughout the sequence, as shown in Fig. 1 (Middle). Further visualization results show that Resformer effectively alleviates attention sinks (Xiao et al., 2023) through avoiding value-state drains (Guo et al., 2024b).
@@
During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2023). Existing $KV$-efficient methods often process keys and values simultaneously. Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer with single layer value (SVFormer). In SVFormer, the queries and keys of all layers share the value from the first layer. Consequently, SVFormer can save nearly half of the $KV$ cache. Additionally, SVFormer is orthogonal to the classical method GQA, and they can be used concurrently. Experimental results show that SVFormer trains significantly faster than the vanilla Transformer. Its performance mainly depends on factors like training sequence length and cumulative learning rate.
@@
We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023, Dubey et al., 2024, Kaplan et al., 2020). We compare different models by their relative training curves against the vanilla Transformer. Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO across all settings. Furthermore, SVFormer trains faster than the vanilla Transformer and performs better when the sequence length is longer.
Figure 1: (Left) Illustration of the relative training loss (loss of target model - loss of vanilla Transformer) curve between different Transformer variants; model size is fixed to be 82M. (Middle) The average entropy of token importance across layers in ResFormer vs. the vanilla Transformer, where token importance is derived from the attention matrix. Lower entropy indicates more focused attention on specific tokens. More details can be found in Eqn. 11. (Right) The average entropy of token importance across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023).
@@
Figure 2: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention. $A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value states, and attention outputs at the $i$-th layer, respectively. $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
@@ 2.1 Shortcut Connections For Better Information Flow
Related research indicates that for advanced Transformer architectures, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024). Furthermore, Zhou et al. (2021) illustrates that a 32-layer ViT underperforms a 24-layer ViT. Depth-Wise Attention (ElNokrashy et al., 2024) allows each query to access the key and value at the same position from previous layers through an attention-like mechanism before the output layer. DenseFormer (Pagliardini et al., 2024) integrates weighted fusion of outputs from all preceding layers after each layer. To explore why increasing depth in Transformers does not yield expected gains, Wang et al. (2022) finds that self-attention acts as a low-pass filter, smoothing token representations in ViTs. Additionally, Shi et al. (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks. NeuTRENO (Nguyen et al., 2023) adds the difference between the value states of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
@@ 2.2 $KV$ cache compressing
For non-Transformer methods, Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023) are two popular works. They replace the original softmax-based attention with SSM (Gu et al., 2021) and AFT (Zhai et al., 2021) mechanisms, respectively. Besides, several approaches have been proposed to enhance models’ ability to process long texts while reducing the reliance on $KV$ cache. Dai (2019) advocates segmenting long texts into smaller parts for attention computation. Building upon this, Munkhdalai et al. (2024) proposes using a fixed-size memory matrix for storing and retrieving historical information.
Transformer-based methods can be categorized into three main groups. The first group consists of post-training methods like SnapKV (Li et al., 2024) and ThinK (Xu et al., 2024), which compress KV cache during inference based on attention matrices at token or hidden dimension levels. The second group focuses on quantization and adopts low-precision $KV$ cache quantization rather than completely eliminating them (Hooper et al., 2024). The third group aims to maximize the efficiency of attention-based models via parameter or activation value sharing. The most representative works include Multi-Query Attention (Shazeer, 2019) and Grouped-Query Attention (Ainslie et al., 2023) which suggest to share key and value across a group of queries. MLKV (Zuhri et al., 2024) further suggest to share keys and values for queries across layers and MLA (Liu et al., 2024) introduces low-rank projection when processing keys and values. Besides, CLA (Brandon et al., 2024) and LISA (Mu et al., 2024) respectively point out that we can reuse keys, values, or the attention matrix across layers to reduce redundancy between layers. While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key. It shares values across all layers and is compatible with other methods like GQA.
@@ 3.2 Efficient Cross Layer Attention
Figure 3: Average token similarity between the outputs of different mapping methods and that of Eqn. 2.
Figure 4: Ablation study on sharing keys or values in every two layers, with CLAttention denoting sharing both.
where $\mathbf{A}_{n}$ denotes the original attention matrix for layer $n$. From the perspective of information propagation, model described by Eqn. 3 projects the historical values into the current layer’s embedding space using the current layer’s attention as a weight matrix. For example, a naive approach would be to perform identity mapping, as described by
$\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6)
To evaluate the approximation effect of replacing the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$, we randomly select 1,000 pre-training data samples. For each layer of a trained baseline model, assuming cross-layer attention defined by Eqn. 2 is required for each layer with respect to the previous one, we calculate the cosine similarity between the outputs from Eqn. 2 and Eqn. 3. We also calculate the cosine similarity between the outputs from Eqn. 2 and Eqn. 6 for comparison. Fig. 4 shows that using current attention as a mapping matrix provides a better approximation for cross-layer attention.
@@ 3.4 A unified View Of NeuTRENO and DenseFormer
$\mathbf{U}_{n}$ $=\Big{(}\mathbf{A}_{n}-\bm{\lambda}\mathbf{J}\Big{)}{\mathbf{V}_{n}}+\bm{\lambda}{\mathbf{V}_{1}}.$ (8)
where $\mathbf{J}$ denotes a matrix of ones and $\bm{\lambda}$ is a hyper-parameter. It can be found that the term of $\bm{\lambda}\mathbf{J}$ may have certain negative impact on the learning of original attention. If we ignore the attention output projections and the MLP layers, DenseFormer can also be modeled within our framework as
@@
where $\{\bm{\alpha}_{i}\}_{i=1}^{n}$ is a set of hyper-parameters. DenseFormer uses attention matrix of previous layer as the weight matrix of projecting values but this is not aligned with the concept of the efficient cross layer attention shown in Eqn. 3.
@@ 3.5 SVFormer: Single-layer Value For Half $KV$ Cache
After ResFormer, a natural idea is whether we can remove the value states in each layer and have all layers share the value states from the first layer. We call this method SVFormer. Similar to ResFormer, SVFormer still adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through
@@
Compared to previous methods, SVFormer is the first method that decouple value states from attention. Its main advantage is that it only requires computing and storing the value states for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value states every two layers, but the results in Fig. 4 show that sharing values has less negative impact on performance compared with sharing keys.
@@ 4.1.1 Training Details
Unless otherwise noted, we train all models using AdamW optimizer with 0.1 weight decay, $\beta_{1}=0.9$, $\beta_{2}=0.95$ and the max grad norm 1.0. The batch size is set to be around 2M tokens (Zhang et al., 2024) with a sequence length of 2,048 and the total steps is fixed 10,000 steps (Kaplan et al., 2020). We adopt linear learning rate warmup for the first 1,200 steps with the initial learning rate and the peak learning rate to be 1e-7 and 6e-4 respectively. The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Brandon et al., 2024, Zhou et al., 2024, Wei et al., 2023). See Table 4 and Table 3 in Appendix for more details.
All models are trained with 8 Nvidia A100 80G GPUs using mixed-precision training in FP16. We adopt deepspeed zero-2 optimizer and flash attention mechanism.
Figure 5: (Left) The relative training curve between a 82M ResFormer and Transformer across different training sequence lengths. (Middle) Average training loss for the final 50 steps across different model sizes and the corresponding fitted curves. (Right) The relative training curve across different model size for a fixed 2,048 training sequence length.

### v3

@@ Abstract
Transformers can capture long-range dependencies using self-attention, allowing tokens to attend to all others directly. However, stacking multiple attention layers leads to attention concentration. One natural way to address this issue is to use cross-layer attention, allowing information from earlier layers to be directly accessible to later layers. However, this approach is computationally expensive. To address this problem, we propose Transformer with residual value (ResFormer) which approximates cross-layer attention through adding a residual connection from the values of the the first layer to all subsequent layers. Based on this method, one variant is the Transformer with single layer value (SVFormer), where all layers share the same value embedding from first layer. Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 10.4% fewer model parameters and 13.6% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer reduces KV cache size by nearly half with only a small performance penalty and can be integrated with other KV-efficient methods, yielding further reductions in KV cache, with performance influenced by sequence length and cumulative learning rate. Further visualization results suggest that Resformer and SVFormer alleviate attention concentration in deeper layers through avoiding value-state drains and enhance representation across most layers.  [^1] Code is available at https://github.com/Zcchill/Value-Residual-Learning.
Figure 1: (Left) The average entropy of token importance and the average hidden-state similarity for a randomly initialized 468M model. (Middle) The average entropy of token importance across layers in Llama (8B) (Dubey et al., 2024) and Mistral (7B) (Jiang et al., 2023). (Right) The average entropy of token importance across layers in ResFormer vs. the vanilla Transformer, where token importance is derived from the attention matrix. Lower entropy indicates more focused attention on specific tokens. More details can be found in Eqn. 11.
@@ 1 Introduction
The Transformer (Vaswani et al., 2017) model has become one of the leading architectures in recent years, excelling in both language modeling (Devlin et al., 2019, Lan et al., 2020, Brown et al., 2020) and computer vision tasks (Dosovitskiy et al., 2021). The discovery of scaling laws (Hoffmann et al., 2022, Kaplan et al., 2020) has driven the pursuit of larger Transformer models by increasing network depth and width.
(a) Token importance.
(b) Norms of value states.
(c) Norms of hidden states.
Figure 2: The token importance (Xiao et al., 2024), value-state norms (Guo et al., 2024b), and hidden-state norms (Sun et al., 2024) of the first token across layers of 468M models. More Visualization results are available in Appendix A.2.
@@
Existing solutions to alleviate the over-smoothing problem in Transformer include adding extra regularizers (Nguyen et al., 2023, Shi et al., 2022) and optimizing the information flow within the model (Pagliardini et al., 2024). During the era of convolutional neural network architectures, Stochastic Depth (Huang et al., 2016) reduces the likelihood of over-smoothing by randomly dropping layers during training and DenseNet (Huang et al., 2017) mitigates the impact of over-smoothing by allowing each layer to directly access the hidden states of all preceding layers. Recently, DenseFormer (Pagliardini et al., 2024) adopts the idea of DenseNet when training Transformer. Additionally, NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing through incorporating the difference between the value vectors of the first layer and the current layer to the attention output.
@@
In this paper, we address the problem of multi-layer attention from another perspective. We introduce the phenomenon of attention concentration, which describes how a model’s attention increasingly focuses on fewer tokens. We quantify the degree of attention concentration using the entropy of the distribution of token importance, where lower entropy indicates a more pronounced concentration. Unlike over-smoothing, which is inherent to model architecture, attention concentration emerges during training. Fig. 1 (Left) shows that randomly initialized models exhibit over-smoothing but not attention concentration. Trained ViT models often focus on low-informative background areas (Darcet et al., 2024), while language models concentrate on low-semantic tokens (Sun et al., 2024), particularly the start token (attention sink (Xiao et al., 2024)). While previous studies analyzed single-layer attention patterns, our research reveals a “concentration - dispersion - concentration” pattern in deep models, as shown in Fig. 1 (Middle), suggesting potential loss of information during concentrated phases. The analysis of over-smoothing is available in Appendix A.1.
Mitigating attention concentration can lead to more interpretable attention maps and potentially improve downstream task performance (Darcet et al., 2024). This phenomenon typically emerges after the second or third network layer and is associated with value-state drains (decreased magnitude of value states) (Guo et al., 2024b), and hidden-state peaks (increased magnitude of hidden states) (Sun et al., 2024). Guo et al. (2024a) shows a mutual reinforcement mechanism exists between value-state drains and attention concentration. Recent studies have linked this to implicit biases during pretraining, with most existing solutions focusing on the use of additional tokens (registers) (Darcet et al., 2024) or additional keys and values (explicit attention bias) (Sun et al., 2024) to redirect this.
Given that the first layer always shows no attention concentration, an effective method is to use cross-layer attention on information from this layer. However, due to computational costs, we propose ResFormer as an efficient alternative. ResFormer applies a residual connection between the value vectors of the current layer and the first layer before the attention operation. Unlike cross-layer attention, ResFormer indirectly mitigates attention concentration. It leverages the absence of value-state drains in the first layer by introducing a value residual connection. This alleviates value-state drains in deeper layers, thereby disrupting the mutual reinforcement between attention concentration and value-state drains, as shown in Fig. 1 (Right) and Fig. 2.
@@
During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2024). Existing $KV$-efficient methods often process keys and values simultaneously. Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer with single layer value (SVFormer). In SVFormer, the queries and keys of all layers share the value from the first layer, and thus it can also alleviate attention concentration.
@@
We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023, Dubey et al., 2024, Kaplan et al., 2020). We compare different models by their relative training curves against the vanilla Transformer. Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO. ResFormer achieves equivalent validation loss with 10.4% fewer model parameters and 13.6% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer. And the performance of SVFormer is better when the training sequence length is longer. It further reduces the $KV$ cache when integrated with classical method GQA (Ainslie et al., 2023).
@@
Figure 3: Simplified illustration of the vanilla Transformer, NeuTRENO, DenseFormer, ResFormer, and SVFormer, with only three-layer structures and no operations other than attention. $A^{i}$, $V^{i}$, and $H^{i}$ denote the attention matrix, value vectors, and attention outputs at the $i$-th layer, respectively. $\oplus$, $\ominus$, and $\otimes$ represent standard matrix addition, subtraction, and multiplication, respectively.
@@ 2.1 Shortcut Connections For Better Information Flow
Related research indicates that for advanced Transformer architectures, although increasing depth continues to yield performance improvements in language modeling tasks, the gains become less significant with further increases (Petty et al., 2024). Furthermore, Zhou et al. (2021) illustrates that a 32-layer ViT underperforms a 24-layer ViT. Depth-Wise Attention (ElNokrashy et al., 2024) allows each query to access the key and value at the same position from previous layers through an attention-like mechanism before the output layer. DenseFormer (Pagliardini et al., 2024) integrates weighted fusion of outputs from all preceding layers after each layer. To explore why increasing depth in Transformers does not yield expected gains, Wang et al. (2022) finds that self-attention acts as a low-pass filter, smoothing token representations in ViTs. Additionally, Shi et al. (2022) investigates over-smoothing from a graph perspective in BERT-based language modeling tasks. NeuTRENO (Nguyen et al., 2023) adds the difference between the value vectors of the first and current layers to each layer’s attention output and significantly alleviates the over-smoothing problem.
@@ 2.2 $KV$ cache compressing
For non-Transformer methods, Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023) are two popular works. They replace the original softmax-based attention with SSM (Gu et al., 2021) and AFT (Zhai et al., 2021) mechanisms, respectively. Besides, several approaches have been proposed to enhance models’ ability to process long texts while reducing the reliance on $KV$ cache. Dai et al. (2019) advocates segmenting long texts into smaller parts for attention computation. Furthermore, Munkhdalai et al. (2024) uses a fixed-size memory matrix for storing and retrieving past information.
Transformer-based methods can be categorized into three main groups. The first group consists of post-training methods like SnapKV (Li et al., 2024) and ThinK (Xu et al., 2024), which compress $KV$ cache during inference based on attention matrices at token or hidden dimension levels. The second group focuses on quantization and adopts low-precision $KV$ cache quantization rather than completely eliminating them (Hooper et al., 2024). The third group aims to maximize the efficiency of attention-based models via parameter or activation value sharing. The most representative works include Multi-Query Attention (Shazeer, 2019) and Grouped-Query Attention (Ainslie et al., 2023) which suggest to share key and value across a group of queries. MLKV (Zuhri et al., 2024) further suggest to share keys and values for queries across layers and MLA (Liu et al., 2024) introduces low-rank projection when processing keys and values. Besides, CLA (Brandon et al., 2024) and LISA (Mu et al., 2024) respectively point out that we can reuse keys, values, or the attention matrix across layers to reduce redundancy between layers. While these methods typically process both key and value simultaneously, SVFormer is the first approach to decouple value from query and key during attention computation. Moreover, it is compatible with other methods like GQA.
@@ 3.2 Efficient Cross Layer Attention
where $\mathbf{A}_{n}$ denotes the original attention matrix for layer $n$. From the perspective of information propagation, model described by Eqn. 3 projects the historical values into the current layer’s embedding space using the current layer’s attention as a weight matrix. For example, a naive approach would be to perform identity mapping, as described by $\mathbf{U}_{n}$ $=\mathbf{A}_{n}\mathbf{V}_{n}+\frac{1}{n-1}\sum_{i=1}^{n-1}{\mathbf{V}_{i}}.$ (6) To evaluate the approximation effect of replacing the $\mathbf{K}_{m}$ with $\mathbf{K}_{n}$, we randomly select 1,000 pre-training data samples. For each layer of a trained baseline model, assuming cross-layer attention is required for each layer with respect to the previous one, we calculate the cosine similarity between the outputs from Eqn. 2 and Eqn. 4. We also calculate the cosine similarity between the outputs from Eqn. 2 and Eqn. 6 for comparison. Fig. 4 shows that our proposed method provides a good approximation for cross-layer attention.     [Uncaptioned image] Figure 4: Average token similarity between layer outputs. Lines show similarity of outputs using current attention (Eqn. 4) or identity attention (Eqn. 6) compared to the one using cross-layer attention (Eqn. 2).
@@ 3.4 A unified View Of NeuTRENO and DenseFormer
$\mathbf{U}_{n}$ $=\Big{(}\mathbf{A}_{n}-\bm{\lambda}\mathbf{I}\Big{)}{\mathbf{V}_{n}}+\bm{\lambda}{\mathbf{V}_{1}}.$ (8)
where $\mathbf{I}$ denotes the identity matrix and $\bm{\lambda}$ is a hyper-parameter. It can be found that the term of $\bm{\lambda}\mathbf{I}$ may have certain negative impact on the learning of original attention. If we ignore the attention output projection and the MLP layer, DenseFormer can also be modeled within our framework as
@@
where $\{\bm{\alpha}_{i}\}_{i=1}^{n}$ is a set of hyper-parameters. DenseFormer uses attention matrix of previous layer as the weight matrix of projecting values but this is not aligned with the concept shown in Eqn. 3.
@@ 3.5 SVFormer: Single-layer Value For Half $KV$ Cache
After ResFormer, a natural idea is whether we can remove the value vectors in each layer and have all layers share the value vectors from the first layer. We call this method SVFormer. Similar to ResFormer, SVFormer still adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through
@@
Compared to previous methods, SVFormer is the first method that decouple value vectors from attention. Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers. However, the results in Fig. 5 show that sharing values has less negative impact compared with sharing keys.
Figure 5: Ablation study of sharing different parts of attention every two layers.
@@ 4.1.1 Training Details
Unless otherwise noted, we train all models using AdamW optimizer with 0.1 weight decay, $\beta_{1}=0.9$, $\beta_{2}=0.95$ and the max grad norm 1.0. The batch size is set to be around 2M tokens (Zhang et al., 2024) with a sequence length of 2,048 and the total steps is fixed 10,000 steps (Kaplan et al., 2020). We adopt linear learning rate warmup for the first 1,200 steps with the initial learning rate and the peak learning rate to be 1e-7 and 6e-4 respectively. The cosine decay schedule gradually decays to 10% of the peak learning rate by the end of training (Zhou et al., 2024, Wei et al., 2023). The detailed hyperparameters for models of various sizes and different training sequence lengths in Appendix A.5. Moreover, All models are trained with 8 Nvidia A100 80G GPUs using mixed-precision training in FP16. We adopt deepspeed zero-2 optimizer and flash attention mechanism.